What is the pH of a solution that is 0.18 M K2HPO4 and 0.059 M K3PO4? Use the Acid-Base Table. Please see Buffer pH for assistance. What is the pH of a solution that is 0.18 M K2HPO4 and 0.059 M K3PO4? Use the Acid-Base Table.
Answer #1

mixture of K2HPO4 and K3PO4 act as acidic buffer

pH = pKa + log [K3PO4] / [K2HPO4]

pKa of K2HPO4 = 12.38

[K3PO4] = 0.059 M

[K2HPO4] = 0.18 M

pH = 12.38 + log [0.059] /[0.18]

pH = 12.38 - 0.48

pH = 11.90